20180622 UT - another attempt at AGw3 alignment

Observers: JHill, OKuhn, BRothberg (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: DMiller (Tucson)
Telescope Support: PHartley (LBTO)
Software Support:
Instrument Support: JMorris (LBTO)
Engineering Support: DCox

Summary

We worked from twilight until 04:56 UT on AGw3 alignment.

We refined the AGW3 guiding hotspot (it was already close).

We made a huge correction (15 arcsec) to the probe centering, but have not yet taken the confirming/refining measurement. Take note that the X-flip changes the apparent direction you need to move.

We used a single Stone field mask to measure and confirm the LEFTZEROPOINT.

The next step would be to confirm the probe centering, and then take some transform data. Doug also needs some Z9/Z10 Zernike data to finish active optics, but we could live without that for a while.

Preparation

All this depends on AGW3 being suitably dried out after the glycol leak.....which it was. See IT 7282 for an update on dark current from non-functional TECs. Petr has deployed a patch for GCSGUI to allow X-flip in the guider display.

Plan

DONE Normal pointing correction, then ACTIVE preset in twilight. Collimate with IDL.

DONE Refine guidecam_hotspot_x,_y (BGTechProcedures) [0.2 hour in twilight]

(partly done) Set Probe center position in AGW3 (BGTechProcedures)[0.5 hour, prefer good seeing]

DONE Iterate guidecam_hotspot and WFScam_hotspot with collimation.

Skip Pupil Wobble measurement as SX M3 mounting is unchanged, but beware IT 7269. (Commissioning/ComTools#PupilWobble)

DONE Check LEFTZEROPOINT wrt/LUCI slit or mask [0.5 hour] - MOS masks StoneJ or StoneL

Check location of rotator center on LUCI1. [0.3 hour]

Collect Z9, Z10 collimation data on-axis to refine Zernike rotation angle. [0.5 hour]

Check off-axis guiding and collimation. [0.2 hour] can be part of the transform data set

Collect initial LUCI1 and LUCI2 transform data [1-2 hours, requires good seeing] (We are not anticipating data reduction during the night.)

Details

Open to a beautiful clear sky......Authorized as LUCI/LUCI.

03:00 SX IE=-17 CA=40 DX IE=-28 CA=-42

03:03 Steve is collimating SX with IDL.

03:08 Start adjusting hotspot.

03:10 279,242 0006 wrong sign

03:13 279,252 0007 keep this.

Adjusting probe to out new hotspot on rotator center

Stop guiding

setxy -u 3 -x 0 -y 612.5

03:19 guide 0001 hung

03:23 selectAGW

03:24 guide 0003 center of circle is at 58, 160.5 binned 116,321 detxy, 131,360 native (compare hotspot at 279,252)
decrease rot homeoffset to increase X want +148 pixels = 7.4 mm * 68 =-503 counts
increase rad oofset to decrease Y by 108 pixels = 5400 microns
John forgot to convert from arcsec to mm!
editted oacontrol.conf
rdwrconfig -u 3

03:42 restart GCSL and resend preset (but acquire image is all black) move to brighter star

New SX IE=-5 CA=+57 DX IE=-37 CA=-42

03:53 ACTIVE preset - seeing is 0.6 FWHM at EL=80

03:58 ACTIVE preset to star at EL=68

04:00 guide 0004 but John went the wrong way because he forgot about the X-flip!
now increase rot homeoffset to decrease X want +148 pixels = 7.4 mm * 68 =+503 counts were applied ---> 130945
John forgot to convert from arcsec to mm!
should have said increase rot homeoffset to increase X want +148 pixels = 7.4 arcsec = 4.44 mm * 68 =+302 counts

04:05 rdwrconfig -u 3 restart GCSL THIS CHANGE IS NOT YET VERIFIED. That will be the next step on-sky to take a refining measurement.

04:08 change authorization to LUCI1 only since LUCI2 doesn't have Stone Mask

Measure rotator zeropoint with Stone L mask

04:12 ACQUIRE preset to correct pointing after John moved probe back SX IE=-15 CA=+41
  • Acquisition field images: 14(sky), 15(obj) and Mask image: 16.
  • RTD alignment gives dtheta = +0.11256 deg. (dx,dy = -4.68,-3.83)

04:46 subtract 0.112 deg to LEFTZEROPOINT was -187.805 now -187.917 which was correct, but we'll admit we accidentally got the sign right.
  • Acquisition field images : 18(sky), 19(obj) and Mask image: 20.
  • RTD alignment gives dtheta = -0.012 deg.

04:56 Over to LINC MCAO observing

05:00 Took some empty 30 sec WFS images starting with /lbt/data/share/tcs/lbto.20180622/left_wfsc000032.fits

05:20 Took some empty 15 sec WFS images starting with /lbt/data/share/tcs/lbto.20180622/left_wfsc000040.fits

-- %USERSIG{JohnHill - 2018-06-21)
Topic revision: r5 - 26 Jun 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback